OpenStudy (anonymous):

f(x)=2x-7 and g(x)=-6x-3 find (f+g)(x)

OpenStudy (anonymous):

(f+g)(x)= -4x-10

OpenStudy (jordanloveangel):

f(x)-g(x) = (2x2 -2x + 7) - (x2 + 9x - 6) = (2x2 - x2) + (-2x - 9x) + (7 + 6) = x2 - ... (i think not sure thou )